OpenStudy (anonymous):

A room is 19.4 ft deep, 25.4 ft wide and 7.39 ft high. Find the volume of this room. 1 inch = 2.54 cm. Answer in units of m^3

hartnn (hartnn):

u know how many inches are there in a feet?

OpenStudy (anonymous):

12 obviously

hartnn (hartnn):

yes, then first convert 19.4 feet into inches and tell me what u get.

OpenStudy (anonymous):

I've tried this problem before so here's the conversions I got for all of them: 232.8, 304.8, and 88.68

hartnn (hartnn):

ok, now 232.8 inches = 232.8 *2.54 cm = 232.8 *2.54 * 0.01 m = ??

hartnn (hartnn):

did u get 5.913m ?

OpenStudy (anonymous):

I figured out what I did wrong the first time. I only moved the decimal place once, not twice.

hartnn (hartnn):

ok,so what volume did u get ?

OpenStudy (anonymous):

I already solved the whole problem and got the right answer. thanks for the help though

hartnn (hartnn):

welcome :)